OpenStudy (anonymous):

Please help factoring! x^2+8x+4y+8=0

OpenStudy (anonymous):

x^2+8x__= 4y+8

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Its a parabola that I need to write in standard form, find vertex, focus,directrix, and axis of symmetry

OpenStudy (anonymous):

(x-h)^2=4p(y-k)

OpenStudy (anonymous):

x^2+8x__= 4y+8 Add 16 both side and complete the square on the left hand

OpenStudy (anonymous):



OpenStudy (anonymous):

so vertex is (4,-6)

OpenStudy (anonymous):

(that should be a 4(y+6)

OpenStudy (anonymous):

Your equation is x^2+8x+4y+8=0. right ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Then you should have x^2 + 8x = -4y - 8

OpenStudy (anonymous):

really? ok then vertex is (4,-2)

OpenStudy (anonymous):

and p=-1

OpenStudy (anonymous):

how do i find the focus,symmetry, and directrix?

OpenStudy (anonymous):

Vertex should be 4,2

OpenStudy (anonymous):

oh thats right, theres a -4...

OpenStudy (anonymous):

can you qualitatively draw this parabola ?

OpenStudy (anonymous):

ummm im not sure...

OpenStudy (anonymous):

i can plot the vertex but i dont know any points on the parabola curve

OpenStudy (anonymous):